# Phenomenological study of the angle between jet axes in heavy-ion collisions

### v1

@@ III Results and Discussions
Figure 1: Normalized distributions of the angle between the standard and WTA axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ of inclusive jets calculated by the Pythia in $p$+$p$ collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ within two $p_{T}$ range $40<p_{T}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and $60<p_{T}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ compared to the ALICE data Acharya et al. (2023). The ratio of these two $p_{T}$ ranges is also plotted in the lower panel.
@@
To calibrate the $p$+$p$ baseline of our study, in Fig. 1, we firstly show the normalized distributions of the angle between the standard and WTA axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ of inclusive jets calculated by the Pythia in $p$+$p$ collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ within two $p_{T}$ range $40<p_{T}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and $60<p_{T}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ compared to the ALICE data. It is found that the calculations of Pythia can describe the ALICE data for two $p_{T}$ ranges. By taking the ratio of these two $p_{T}$ ranges in the lower panel, we find that jets with $60<p_{T}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ have narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than that with $40<p_{T}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$, which indicates that higher-$p_{T}$ jets usually have narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ than the lower-$p_{T}$ one.
@@
Now we proceed to the calculations of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in nucleus-nucleus collisions. In Fig. 2, we show the normalized $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02~{}\mathrm{TeV}$ compared to the recent reported ALICE data. At the same time, the ratios of PbPb/$pp$ are also shown in the lower panel. Our theoretical results can reasonably describe the experimental data in both $p$+$p$ and Pb+Pb collisions and the PbPb/$pp$ ratio. We observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb collisions is narrower than that in $p$+$p$  which indicates the two jet axes are closer for the jet sample in Pb+Pb compared to that in $p$+$p$.
@@
Figure 2: Normalized $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distributions for inclusive jets in $p$+$p$ and Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02~{}\mathrm{TeV}$, including the results of theoretical calculation and ALICE experimental data Acharya et al. (2023). The ratios are also plotted in the lower panel.
@@
Generally, the same minimum and maximum $p_{T}$ thresholds (e.g. $40<p_{T}^{\rm jet}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) are imposed to select the jet samples both in $p$+$p$ and A+A collisions. The comparison of these two jet samples provides valuable information about the in-medium jet modification relative to the vacuum case. However, if only considering the jet-medium interaction but not the “selection biases” in the theoretical calculations and experimental measurements, we cannot fully understand the relationship between the energy loss mechanisms and the observables Renk (2013). Consider a jet (unquenched jet) generated in the initial hard scattering, passing through the QGP medium and losing energy due to interaction with the medium, becoming a medium-modified jet (quenched jet). According to the transverse momentum values before and after the energy loss of this jet, we can place this jet in the lower triangle shadow area in Fig. 3. In usual experimental and phenomenological studies of the current, the selected sample of $p$+$p$ events, which is the unquenched jets sample, is located in the I$+$II$+$III region in Fig. 3. The selected jet sample in heavy-ion collisions, the quenched jets, is in the I$+$IV region in Fig. 3. This means the traditional selection method of unquenched and quenched jet samples is significantly biased. We note that the area I corresponds to the jets survived in the selection window (40–60 $\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) after in-medium energy loss. Moreover, area IV corresponds to jets with higher $p_{T}$ ($>60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) but finally falls in the selection window due to the evident energy loss.
@@
The selection bias effect often plays an important role in the studies of nuclear modification in heavy-ion collisions. One can find the related investigations in Refs. Renk (2013); Brewer et al. (2021a); Casalderrey-Solana et al. (2019); Brewer et al. (2021b); Du et al. (2020); Brewer et al. (2022); Cunqueiro and Sickles (2022). For dijet events, some schemes to reduce selection bias have been invented Brewer et al. (2019, 2021b); Du et al. (2020); Takacs and Tywoniuk (2021). In addition, some studies use boson-tagged jet events to reduce the impact of selection bias on the measurement of the nuclear modification for the observable Brewer et al. (2022); Wang et al. (2022). Although it is still challenging to eliminate the influence of selection bias in experimental measurements, one can efficiently track the jet evolution in the Monte Carlo simulations for the theoretical models. In this work, we study the influence of selection bias on our concerned observable from the effect of jet modification using a “matching-jet” method, first proposed in Refs. Brewer et al. (2021b, 2022). Specifically, for the inclusive jet sample, we pair two jets whose distance in the $(\eta,\phi)$ plane is less than the jet radius $R$ before and after quenching, considering them as the unquenched and quenched versions of the same jet. To minimize the omission of jets that have experienced sizeable energy loss, we have appropriately reduced the transverse momentum threshold in the reconstruction of quenched jets ($p_{T}>10~{}\mathrm{GeV}$) to take into account the region of I and II in Fig. 3. Since the contribution of region III is small, and the matching efficiency in this region is poor, we neglect this region in the matching procedure. The matching procedure lets us focus on the jet-by-jet comparison before and after the in-medium evolution. Although this method makes it difficult to consider the effects of the nuclear parton distribution function and could not be applied to experimental measurements, it gives us a new perspective to study the medium modification of the angle between the standard and WTA jet axes.
@@
In Figure 4, we show the distributions of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ for samples of matched jets in the region I$+$II. For comparison, the Monte Carlo simulation results shown in Figure 2 have also been added to this figure. For the convenience of the following discussion, we labeled these matched jets by Select-Then-Quench (STQ), a matching-jet method first introduced in Ref. Brewer et al. (2021b). Since the jets are well-tracked in the matching procedure, comparing the STQ ($pp$) and STQ (PbPb) cases reveals the actual jet modification in the QGP, in which the influences of selection bias are almost removed. We observe that $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of STQ (PbPb) shows evident broadening compared to that of STQ ($pp$), and the PbPb/$pp$ ratio of STQ has enhancement at $\Delta R_{\mathrm{axis}}>0.04$. It means the jets get broader by the interactions with the QGP medium compared to their initial structures. However, the modification patterns of STQ results are inverse to the normal theoretical calculation, while the latter contains the effect of selection bias. Now, we can conclude that the selection bias covers the real jet-broadening nature and leads to the narrowing modification of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution.
@@
This simple test above demonstrates the importance of selection bias in studying the nuclear modification of the $\Delta R_{\mathrm{axis}}$ distribution. However, please note that although we have excluded selection bias using the STQ method, this also obscures the quark-initiated jets increase effect that existed in experiments or conventional Monte Carlo studies, where the proportion of quark-initiated jets tends to increase due to gluon usually losing more energy. In Monte Carlo studies, we can distinguish whether the jet is a quark- or gluon-jet by comparing the distance between the jet and the initial outgoing hard partons in the $(\eta,\phi)$ plane. Utilizing this method, we have employed the Pythia8 event generator to ascertain that the quark- and gluon-jet fractions of inclusive jets in $p$+$p$ collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ at leading-order (LO) accuracy are about 36.7% and 63.3% with jet radius $R=0.2$, $|\eta_{\mathrm{jet}}|<0.7$ and $40<p_{T}^{\mathrm{ch~{}jet}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$. To test the flavor dependence of jet $\Delta R_{\mathrm{axis}}$ and its medium modification, as shown in Figure 5, we estimate the $\Delta R_{\mathrm{axis}}$ distributions of quark-jets, gluon-jets and inclusive jets in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$. The quark-jets exhibit a narrower initial $\Delta R_{\mathrm{axis}}$ distribution than the gluon-jets. In Pb+Pb collisions, we find that both the quark- and gluon-jets become narrower than in $p$+$p$, and the narrowing of the latter is more substantial. Moreover, the $\Delta R_{\mathrm{axis}}$ modification of inclusive jets seems closer to that of the gluon-jets sample.
@@
To test the influence of the quark/gluon-jet fraction change on the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ modification, we rescale their fractions in Pb+Pb collisions to be consistent with their initial fraction (36.7% and 63.3%). The calculations with rescaled quark/gluon-jet fractions, denoted as Lbt (recover $q/g$), are shown in Fig. 6 and compared to the normal Lbt calculations as well as the ALICE data. We find that the calculations with rescaled fractions do not significantly differ from the normal Lbt calculations. Since the rescaled calculations exclude the influence of the fraction changes, it hints that the decreased gluon-jet fraction is not the main reason that leads to a narrowing $\Delta R_{\mathrm{axis}}$ distribution of inclusive jet in Pb+Pb collisions.
@@
Figure 7: The $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of the selected jet sample with $40<p_{T}<60~{}\mathrm{GeV}$ in Pb+Pb is compared to its initial counterpart, denoted as $pp$(matched), and that selected in $p$+$p$ collisions, the ratios of PbPb/$pp$ and PbPb/$pp$(matched) are also plotted in the lower panel.
@@
To find out the actual jet modification, in Fig. 4, we have compared the selected jet sample with $40<p_{T}<60~{}\mathrm{GeV}$ in $p$+$p$ collisions with its quenched counterpart in Pb+Pb collisions using the matching-jet method. Similarly, we can also trace the initial counterpart of the jet sample with $40<p_{T}<60~{}\mathrm{GeV}$ selected in Pb+Pb collisions. We denote such matched initial counterpart as $pp$(matched) for simplicity. In Fig. 7, the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of the selected jet sample with $40<p_{T}<60~{}\mathrm{GeV}$ in Pb+Pb is compared to its initial counterpart, the $pp$(matched), and the one with $40<p_{T}<60~{}\mathrm{GeV}$ selected in $p$+$p$ collisions, the ratios of PbPb/$pp$ and PbPb/$pp$(matched) are also plotted in the lower panel. We find that $pp$(matched) has a significantly narrower distribution than the $p$+$p$. The jet sample selected in A+A collisions after jet quenching usually has higher initial $p_{T}$, while the one with higher $p_{T}$ usually has narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ as shown in Fig. 1. When taking the ratio of Pb+Pb to their initial counterpart, we can observe a broadening in $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ as shown in the lower panel of Fig. 7. However, if we only compare the jet samples with $40<p_{T}<60~{}\mathrm{GeV}$ selected independently in $p$+$p$ and Pb+Pb collisions, the broadening turns into narrowing. In other words, when taking the ratio of PbPb/$pp$ in experimental measurements, the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline. Such a biased comparison between $p$+$p$ and A+A conceals the actual jet-broadening effect, and it leads to some observation that seems like jet modification but do not represent the nature of the jet quenching. Although the selection bias could not be avoided completely in the experiment, some efforts have been made in this direction that may be helpful to future measurements Renk (2013); Brewer et al. (2019); Wang et al. (2021a); Lai et al. (2022).

### v2

@@ III Results and Discussions
Figure 1: Normalized distributions of the angle between the Standard and WTA axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ of inclusive jets calculated by the Pythia in $p$+$p$ collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ within two $p_{\mathrm{T}}$ ranges: $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and $60<p_{\mathrm{T}}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$, are compared to the ALICE data [19]. The ratio of these two $p_{\mathrm{T}}$ ranges is also plotted in the lower panel.
@@
To calibrate the $p$+$p$ baseline of our study, in Figure 1, we firstly show the normalized distributions of the angle between the Standard and WTA axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ of inclusive jets calculated by the Pythia in $p$+$p$ collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ within two $p_{\mathrm{T}}$ ranges $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ and $60<p_{\mathrm{T}}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ compared to the ALICE data. It is found that the calculations of Pythia can describe the ALICE data for these two $p_{\mathrm{T}}$ ranges. By taking the ratio of these two $p_{\mathrm{T}}$ ranges in the lower panel, we find that jets with $60<p_{\mathrm{T}}<80~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$ have a narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than those with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$, which indicates that higher-$p_{\mathrm{T}}$ jets usually have a narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ than the lower-$p_{\mathrm{T}}$ ones.
@@
Now we proceed to the calculations of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in nucleus-nucleus collisions. In Figure 2, we show the normalized $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$ compared to the recent reported ALICE data [19]. At the same time, the ratios of PbPb/$pp$ are also shown in the lower panel. Our theoretical results can reasonably describe the experimental data in both $p$+$p$ and Pb+Pb collisions and the PbPb/$pp$ ratio. We observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb collisions is narrower than that in $p$+$p$  which indicates the two jet axes are closer for the jet sample in Pb+Pb compared to that in $p$+$p$.
@@
Figure 2: Normalized $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distributions for inclusive jets in $p$+$p$ and Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$, including the results of theoretical calculation and ALICE experimental data [19]. The ratios are also plotted in the lower panel.
@@
Generally, the same minimum and maximum $p_{\mathrm{T}}$ thresholds (e.g. $40<p_{\mathrm{T}}^{\mathrm{jet}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) are imposed to select the jet samples both in $p$+$p$ and A+A collisions. The comparison of these two jet samples provides valuable information about the in-medium jet modification relative to the vacuum case. However, if only considering the jet-medium interaction but not the “selection biases” in the theoretical calculations and experimental measurements, we cannot fully understand the relationship between the energy loss mechanisms and the observables [46]. Consider a jet (unquenched jet) generated in the initial hard scattering, passing through the QGP medium and losing energy due to interaction with the medium, becoming a medium-modified jet (quenched jet). According to the transverse momentum values before and after the energy loss of this jet, we can place this jet in the lower triangle shadow area in Figure 3. In usual experimental and phenomenological studies of the current, the selected sample of $p$+$p$ events, which is the unquenched jets sample, is located in the I$+$II$+$III region in Figure 3. The selected jet sample in heavy-ion collisions, the quenched jets, is in the I$+$IV region in Figure 3. This means the traditional selection method of unquenched and quenched jet samples is significantly biased. We note that the area I corresponds to the jets survived in the selection window (40–60 $\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) after in-medium energy loss. Moreover, area IV corresponds to jets with higher $p_{\mathrm{T}}$ ($>60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$) but finally falls in the selection window due to the evident energy loss.
@@
The selection bias effect often plays an important role in the studies of nuclear modification in heavy-ion collisions. One can find the related investigations in Refs. [84, 85, 46, 86, 87, 88, 89, 90, 6, 34]. We have noticed that there have been a lot of research (e.g. [91, 88, 89, 92, 90, 93, 94, 95, 96, 97]) attempts to eliminate or reduce the impact of selection bias on the studies of nuclear modifications. Some experimentalists have suggested that the issue of selection bias could be resolved by canceling the transverse momentum cut on the jet spectrum [94, 95, 96, 97]. In addition, one can efficiently track the jet evolution in the Monte Carlo simulations for the theoretical models. In this work, we study the influence of selection bias on our concerned observable from the effect of jet modification using a “matching-jet” method, first proposed in Refs. [88, 90]. Specifically, in the two hadron-level inclusive jet samples, quenched and unquenched, we pair two jets whose distance in the $(\eta,\phi)$ plane is less than the jet radius $R$ before and after quenching, considering them as the unquenched and quenched versions of the same jet. To minimize the omission of jets that have experienced sizeable energy loss, we have appropriately reduced the transverse momentum threshold in the reconstruction of quenched jets ($p_{\mathrm{T}}>10~{}\mathrm{GeV}$) to take into account the region of I and II in Figure 3. Since the contribution of region III is small, and the matching efficiency in this region is poor, we neglect this region in the matching procedure. The matching procedure lets us focus on the jet-by-jet comparison before and after the in-medium evolution. Although this method makes it difficult to consider the effects of the nuclear parton distribution function and could not be applied to experimental measurements, it gives us a new perspective to study the medium modification of the angle between the Standard and WTA jet axes.
@@
In Figure 4, we show the distributions of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ for samples of matched jets in the region I$+$II. For comparison, the Monte Carlo simulation results shown in Figure 2 have also been added to this figure. For the convenience of the following discussion, we labeled these matched jets by Select-Then-Quench (STQ), a matching-jet method first introduced in Ref. [88]. Since the jets are well-tracked in the matching procedure, comparing the STQ ($pp$) and STQ (PbPb) cases reveals the actual jet modification in the QGP, in which the influences of selection bias are almost removed. We observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of STQ (PbPb) shows evident broadening compared to that of STQ ($pp$), and the PbPb/$pp$ ratio of STQ is enhanced at $\Delta R_{\mathrm{axis}}>0.04$. It means the jets get broader by the interactions with the QGP medium compared to their initial structures. However, the modification patterns of STQ results are inverse to the normal theoretical calculation, while the latter contains the effect of selection bias. Now, we can conclude that the selection bias obscures the real nature of intra-jet broadening and leads to a narrowed modification of the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution. Many researchers have been focused on searching for medium-induced intra-jet broadening [60, 26], and recent experimentalists have also reported some positive signals [96, 97, 95].
@@
This simple test above demonstrates the importance of selection bias in studying the nuclear modification of the $\Delta R_{\mathrm{axis}}$ distribution. However, please note that although we have excluded selection bias using the STQ method, this also obscures the quark-initiated jets’ fraction increase effect that existed in experiments or conventional Monte Carlo studies, where the proportion of quark-initiated jets tends to increase because gluons usually lose more energy [98, 99, 100]. In Monte Carlo studies, we can distinguish whether the jet is a quark- or gluon-jet by comparing the distance between the jet and the initial outgoing hard partons in the $(\eta,\phi)$ plane. Utilizing this method, we have employed the Pythia8 event generator to ascertain that the quark- and gluon-jet fractions of inclusive jets in $p$+$p$ collisions at $\sqrt{s}=5.02~{}\mathrm{TeV}$ at leading-order (LO) accuracy are about 36.7% and 63.3% with jet radius $R=0.2$, $|\eta_{\mathrm{jet}}|<0.7$ and $40<p_{\mathrm{T}}^{\mathrm{ch~{}jet}}<60~{}\mathrm{GeV}\kern-0.50003pt/\kern-0.20004ptc$. To test the flavor dependence of jet $\Delta R_{\mathrm{axis}}$ and its medium modification, as shown in Figure 5, we estimate the $\Delta R_{\mathrm{axis}}$ distributions of quark-jets, gluon-jets and inclusive jets in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$. The quark-jets exhibit a narrower initial $\Delta R_{\mathrm{axis}}$ distribution than the gluon-jets. In Pb+Pb collisions, we find that both the quark- and gluon-jets become narrower than in $p$+$p$, and the narrowing of the latter is more substantial. Moreover, the $\Delta R_{\mathrm{axis}}$ modification of inclusive jets seems closer to that of the gluon-jets sample.
@@
To test the influence of the quark/gluon-jet fraction change on the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ modification, we rescale their fractions in Pb+Pb collisions to be consistent with their initial fraction (36.7% and 63.3%). The calculations with rescaled quark/gluon-jet fractions, denoted as Lbt (recover $q/g$), are shown in Figure 6 and compared to the normal Lbt calculations as well as the ALICE data. We find that the calculations with rescaled fractions do not significantly differ from the normal Lbt calculations. Since the rescaled calculations exclude the influence of the fraction changes, it hints that the decreased gluon-jet fraction is not the main reason that leads to a narrowing $\Delta R_{\mathrm{axis}}$ distribution of inclusive jet in Pb+Pb collisions.
@@
Figure 7: The $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of the selected jet sample with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ in Pb+Pb is compared to its initial counterpart, denoted as $pp$(matched), and that selected in $p$+$p$ collisions, the ratios of PbPb/$pp$ and PbPb/$pp$(matched) are also plotted in the lower panel.
@@
To find out the actual jet modification, in Figure 4, we have compared the selected jet sample with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ in $p$+$p$ collisions with its quenched counterpart in Pb+Pb collisions using the matching-jet method. Similarly, we can also trace the initial counterpart of the jet sample with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ selected in Pb+Pb collisions. We denote such matched initial counterpart as $pp$(matched) for simplicity. In Figure 7, the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution of the selected jet sample with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ in Pb+Pb is compared to its initial counterpart, the $pp$(matched), and the one with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ selected in $p$+$p$ collisions; the ratios of PbPb/$pp$ and PbPb/$pp$(matched) are also plotted in the lower panel. We find that $pp$(matched) has a significantly narrower distribution than the $p$+$p$. The jet sample selected in A+A collisions after jet quenching usually has higher initial $p_{\mathrm{T}}$, while the one with higher $p_{\mathrm{T}}$ usually has narrower $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ as shown in Figure 1. When taking the ratio of Pb+Pb to their initial counterpart, we can observe a broadening in $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ as shown in the lower panel of Figure 7. However, if we only compare the jet samples with $40<p_{\mathrm{T}}<60~{}\mathrm{GeV}$ selected independently in $p$+$p$ and Pb+Pb collisions, the broadening turns into narrowing. In other words, when taking the ratio of PbPb/$pp$ in the ALICE measurements, the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline. Such a biased comparison between $p$+$p$ and A+A might conceal the real intra-jet broadening behavior during the jet quenching.
Finally, another point to note is that our work has not considered nPDFs effects in the Pb+Pb collision simulations. Our study investigated how cold nuclear effects affect $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distributions in Pb+Pb collisions. The overall trend of nuclear modification showed no significant difference whether we included these cold nuclear effects in the model or not. To facilitate the application of the “matching-jet” method for matching Pb+Pb events with $p$+$p$ references, we deliberately omitted nPDF effects in this work.